### v1

# Monotone Subsystem Decomposition for Efficient Multi-Objective Robot Design

Abstract
Automating design minimizes errors, accelerates the design process, and reduces cost. However, automating robot design is challenging due to recursive constraints, multiple design objectives, and cross-domain design complexity possibly spanning multiple abstraction layers. Here we look at the problem of component selection, a combinatorial optimization problem in which a designer, given a robot model, must select compatible components from an extensive catalog. The goal is to satisfy high-level task specifications while optimally balancing trade-offs between competing design objectives. In this paper, we extend our previous constraint programming approach to multi-objective design problems and propose the novel technique of monotone subsystem decomposition to efficiently compute a Pareto front of solutions for large-scale problems. We prove that subsystems can be optimized for their Pareto fronts and, under certain conditions, these results can be used to determine a globally optimal Pareto front. Furthermore, subsystems serve as an intuitive design abstraction and can be reused across various design problems. Using an example quadcopter design problem, we compare our method to a linear programming approach and demonstrate our method scales better for large catalogs, solving a multi-objective problem of 1025 component combinations in seconds. We then expand the original problem and solve a task-oriented, multi-objective design problem to build a fleet of quadcopters to deliver packages. We compute a Pareto front of solutions in seconds where each solution contains an optimal component-level design and an optimal package delivery schedule for each quadcopter.
I INTRODUCTION
Designing robots, especially optimal ones, is challenging for several reasons. For instance, the search space of possible robot designs is large and can be both continuous and discrete. Moreover, many physics-based robot models have recursive constraints that need to be solved iteratively at each time-step and make optimization difficult. Finally, many approaches are heuristics-based and provide no guarantees of optimality of the final design.
In automated robot design, an often overlooked challenge is the problem of component selection. During this stage, a designer must select components for a predefined robot model to meet high-level task specifications while optimizing one or more objectives. This is a combinatorial optimization problem that remains NP-hard even when formulated as an integer programming problem [1]. If one considers that parts suppliers such as DigiKey or Mouser111www.digikey.com and www.mouser.com have catalogs that include tens of thousands of components, and designs are often optimized for several objectives, it becomes apparent that any approach must use sophisticated techniques to keep the problem tractable.
Figure 1: We use our monotone subsystem decomposition technique to solve a large scale, multi-objective design problem where a fleet of quadcopters must deliver packages. By optimizing subsystems for individual Pareto fronts (PFs) and using these results to optimize the larger system, we can efficiently compute a globally optimal PF of solutions in seconds, where each solution on the PF provides the component-level design for each quadcopter and an individual delivery schedule for the packages that quadcopter has been assigned.
In previous work [2], we presented a framework using constraint programming (CP) for the single-objective component selection problem. In this paper, we extend this framework to multi-objective robot design, which is notoriously difficult due to the large search space. To address this, we introduce a new way to re-use optimization results of subsystems and still guarantee global optimality. This technique speeds up computation by solving smaller problems that have smaller search spaces; saves future computation as optimized subsystems can be reused; and provides a designer useful insight into the optimal design of internal subsystems.
We compare our CP subsystem approach to a linear programming (LP) method from [3] using a quadcopter design problem. Using component catalogs that are several orders of magnitude larger than previous related work, we demonstrate the scaling efficiency of our approach which can solve a multi-objective problem of 1025 candidate solutions and produce a Pareto front (PF) of optimal solutions faster than the LP approach can solve the same single-objective problem.
We further showcase the advantages of our approach by demonstrating its use in a large, complex case study: expanding the previous quadcopter design problem to design a quadcopter fleet for a multi-agent package delivery task, minimizing both delivery time and cost. We can produce a PF of solutions (each consisting of a set of optimal, component-level quadcopter designs and a set of optimal schedules for each quadcopter) in seconds, see Figure 1.
Our primary contributions are as follows:
  • A decomposition technique for constrained optimization problems which allows reusing the optimization results of subsystems under certain consistency conditions
  • Using this subsystem decomposition to solve a combined component selection and task planning problem across four levels of abstraction
In the remainder of this paper, we first summarize related work (Sec. II) and introduce notation for single- and multi-objective constrained optimization problems (COPs) (Sec. III). We then provide the foundational framework for describing monotone subsystem decomposition (Sec. IV). In Sec. V we prove under certain conditions that the PF of a consistent subsystem can be used to determine a PF of globally optimal solutions, which is the theoretical basis for monotone subsystem decomposition. Sec.VI highlights advantages, discusses practical considerations, and contrasts our method with MCDPs, highlighting our method’s greater flexibility. Finally, computational experiments validating our approach are presented in Sec. VII.
II RELATED WORK
Many of the current formalizations and methodologies used in modern robotics are inadequate for automating robot design as they primarily address single-domain challenges [4]. Thus, specific efforts must be made to solve the problem of optimal robot co-design, which requires methods that can cross one or more levels of abstraction.
One approach to the co-design problem uses model-free abstractions in which the robot’s design is not predetermined. Instead, a high-level goal guides the design process to achieve the desired functionality. Model-free approaches include evolutionary algorithms [5, 6], graph heuristic approaches [7], human-assisted design [8, 9, 10], and modular approaches [11]. While these methods lack a higher-level model dictating subcomponent connections, the subcomponents and environment are often well-modeled to facilitate simulation- or heuristics-based optimization. Additionally, they do not guarantee optimal solutions.
In contrast to model-free approaches, model-based approaches define an explicit model for the entire robot design. These approaches are well-suited for the problem of component selection (where components must be selected from large catalogs) as the model can be used to specify component types and relationships between these components. While this restricts the design space, model-based design enables formal guarantees on the system.
The most common model-based methods are linear and quadratic programming. These methods leverage extensive prior research and highly optimized libraries to achieve fast component selection in [12] and [3]. However, designers must linearize the system, which is challenging due to the prevalence of non-linear dynamics.
A second model-based approach is Monotone Co-Design Problems (MCDPs). MCDPs are a framework to describe multi-objective systems where monotone subsystems are connected via functionalities and resources [13]. They can solve multi-objective problems with non-linear, non-convex, or non-continuous systems [14], but require all relationships between functionalities and resources to be monotone. While inspired by MCDPs, our formulation supports a broader range of constraints and focuses on scalability for large component catalogs [2].
The proposed approach and our previous work [2] use constraint programming, also a model-based approach. Constraint programming works well for component selection as it is compatible with non-linear systems and can work with large component catalogs that have discrete component choices ill-suited for linearization. It also provides a more natural representation for the problem of component selection and, as demonstrated in Sec. VII, scales more efficiently than binary integer linear programming.
Our monotone subsystem decomposition technique is most similar to the hierarchical Pareto optimization approach of Singh et al., applied in [15] to design sustainable industrial ecosystems. The advantage of our approach is that we do not need to iterate on solved subsystems – once a subsystem is optimized, it is not necessary to re-optimize. Our subsystem decomposition requires monotonicity, but we also prove that solutions that use subsystem PFs are globally optimal.
III NOTATION
III-A Single-Objective Constrained Optimization
Single-objective COPs are a class of optimization problems that attempt to find a solution that maximizes or minimizes an objective while satisfying a set of constraints. In  [2], we presented a framework for using the constrained optimization representation to solve the problem of component selection, where one knows a general robot model and needs to select specific components from a large catalog. In our formulation, we use a finite index set for variables ${V}=\{1,2,\ldots,{N_{V}}\}$, a set of finite domains (one for each variable index $v$) $\mathcal{D}=\{D_{1},\ldots,D_{N_{V}}\}$, and a finite index set for constraints $C=\{1,2,\ldots,{N_{C}}\}$. $d_{v}\in D_{v}$ are tuples of length $N_{D_{v}}$, i.e. $|d_{v}|=N_{D_{v}}$ where tuple lengths may vary across variable domains but are uniform within each domain. Each element in the tuple $d_{v}$ corresponds to a component property for variable $v$, and these properties are used to characterize the component and describe relationships between other components. We index the variable properties $d_{v}$ using an index set ${P}_{v}=\{1,\ldots,N_{D_{v}}\}$, defined for each $v\in V$, so that the p-th property of a variable v can be written $D_{v,p}$.
Constraints are mathematical expressions that describe relationships between the properties of components. These constraints can not only specify valid component combinations but also describe the performance specifications of the robot. Inequality constraints take the general form of $L_{c}(D_{v,p},D_{v^{\prime},p^{\prime}},...)\leq R_{c}(D_{v,p},D_{v^{\prime},p^{\prime}},...)$ where $L_{c}:\mathbb{R}^{\sum_{v}|P_{v}|}\rightarrow\mathbb{R}$ and $R_{c}:\mathbb{R}^{\sum_{v}|P_{v}|}\rightarrow\mathbb{R}$ are real-valued functions that operate on the variable properties and describe constraint $c\in C$.
The goal is to find an assignment $\mathbf{v}$ that maximizes or minimizes an objective function $f_{i}(\mathbf{v})$ subject to the constraints. Objective functions can also be defined on variable properties $f_{i}(\mathbf{v})\equiv f_{i}(D_{v,p},D_{v^{\prime},p^{\prime}},...):\mathbb{R}^{\sum_{v}|P_{v}|}\rightarrow\mathbb{R}$, and we denote $f_{i}^{\downarrow}$ and $f_{i}^{\uparrow}$ according to whether these objective functions should be minimized or maximized, respectively. Note we use a lowercase $v$ for indexing and a bold $\mathbf{v}$ for assignments. Solving a COP can be formulated as a backtracking tree-search, see our previous work for more details [2].
To aid development of the constraint-based model, it is often easier to use descriptive names, and we associate finite strings with elements of the different index sets. Let the mappings $h_{V}:{V}\rightarrow\Sigma^{*}$, $h_{{{P}_{v}}}:{P}_{v}\rightarrow\Sigma^{*}$, and $h_{{C}}:C\rightarrow\Sigma^{*}$ associate unique names (finite strings over a finite alphabet $\Sigma$) to variables, their properties, and constraints, improving readability when writing equations. For example, in our quadcopter problem, a battery variable represents different battery choices and has properties such as current and voltage. Using these mappings, we write $D_{v,p}$ as $B_{\textrm{voltage}}$, where $D_{v,p}=D_{h^{-1}_{V}(B),h^{-1}_{P_{v}}(\textrm{voltage})}$ and $h^{-1}_{V}$ and $h^{-1}_{P_{v}}$ are the inverse mappings of $h_{V}(v)=B$ and $h_{P_{v}}(p)=\textrm{voltage}$. This shorthand allows constraints to be conveniently expressed as:
$4*(M_{\textrm{voltage}}*M_{\textrm{current}})\vspace{-1.25mm}\leq B_{\textrm{voltage}}*B_{\textrm{current}}$
where the terms of the left and right hand side refer to the associated voltage and current properties of the motor (M) and battery (B) variables, respectively.
III-B Multi-Objective Constrained Optimization
Multi-objective optimization aims to simultaneously optimize several objective functions to obtain an efficient front of solutions (i.e., a PF). A vector of objective functions ${F}(\mathbf{v})=(f_{1}(\mathbf{v}),...,f_{N_{F}}(\mathbf{v}))$ represents the quantities to be optimized, with a corresponding index set $I=\{1,\ldots,N_{F}\}$. This induces a partial order on the search space, where (assuming minimization $\forall f_{i}\in F$) an assignment $\mathbf{a}$ “dominates” another assignment $\mathbf{b}$, i.e., $\mathbf{a}\preceq\mathbf{b}$, if $f_{i}(\mathbf{a})\leq f_{i}(\mathbf{b}),\forall i\in I$ and at least one inequality is strict. The optimal PF $O_{V}$ is the set of non-dominated solutions $\{\textbf{v}\in\textbf{V}\ |\ \nexists\textbf{v}^{\prime}\in\textbf{V},\textbf{v}^{\prime}\preceq\textbf{v}\}$.
To solve a multi-objective optimization problem, a solver maintains an upper bound frontier and computes a lower bound frontier, backtracking if the upper bound dominates the lower. More advanced methods (e.g. AND/OR branch-and-bound [16] or Russian doll search [17]) efficiently explore the search space. However, widely available industrial solvers such as IBM CPLEX [18] and Gurobi [19] do not perform true multi-objective optimization (instead using lexicographical or weighted objectives).
In this paper, we use a different technique to determine the PF of solutions, choosing to take advantage of these highly optimized libraries (in particular IBM CPLEX). We use $k$th-objective $\epsilon$-constraint scalarization [20], where one objective $f_{k}$ is minimized while the rest are constrained by upper bounds $f_{i}(x)<\epsilon,\forall i\in I,i\neq k$. In our case, we choose upper bounds of previously found solutions and perform a lexicographical single-objective solve, adding a constraint to ensure new solutions improve at least one objective (i.e. a new solution must not be dominated by the current PF). We iteratively solve for new solutions, adding new constraints until no new solutions are found. The resulting set of optimal solutions forms a PF.
IV PROBLEM FORMULATION
Using the notation from Sec. III, we now describe the foundational framework for describing monotone subsystem decomposition.
Figure 2: A constraint graph for the Motor and ESC subsystems in Sec. VII. The voltage regulator (VR), half bridge (HB), microcontroller (MC), motor (M), and propeller (P) components are represented as circles, and constraint and objective functions as squares. An edge connects a variable to a function if that variable has a property in the domain of the function; select edges and their properties have been labeled. A subsystem (blue dashed line) is a subset of component nodes. If fully consistent, the subsystem’s PF can be used to optimize the larger system.
To keep track of constraint and objective function dependencies on variables, we use a bipartite multi-graph. One set of nodes represents variables and the other set represents functions of the properties, see Fig. 2. An edge between a function and a variable indicates that the function depends on a property of that variable. Parallel edges indicate a function has multiple properties of the same variable in its domain.
When functions are consistently increasing or decreasing, subsystems can be optimized independently and their results used to optimize the larger system. A variable property is monotone if increasing its value never decreases the function value, and antitone if increasing its value never increases the function value. A variable property is consistent if it is either a consistent maximization property $D^{\uparrow}_{v,p}$ or a consistent minimization property $D^{\downarrow}_{v,p}$. A variable property is a consistent maximization property $D^{\uparrow}_{v,p}$ iff it appears exclusively in inequality constraints where it is antitone for $L_{c}$ and $f_{i}^{\downarrow}$ and monotone for $R_{c}$ and $f_{i}^{\uparrow}$. Likewise, a variable property is a consistent minimization property $D^{\downarrow}_{v,p}$ iff it appears exclusively in inequality constraints where it is monotone for $L_{c}$ and $f_{i}^{\downarrow}$ and antitone for $R_{c}$ and $f_{i}^{\uparrow}$. A constraint function $L_{c}$, $R_{c}$, or objective function $f_{i}$ is fully consistent if all variable properties in its domain are consistent.
We can determine a subsystem ${X}$ within the constraint graph by selecting a subset of variables from the larger system ${Z}={X}\cup{Y}$, where ${X}\subset{Z}\subseteq{V}$ and ${Y}={Z}\setminus{X}$. We denote an implementation $\mathbf{z}$ of subsystem ${Z}$ that uses implementations $\mathbf{x}$ and $\mathbf{y}$ from ${X},{Y}$ as $\mathbf{z}=\mathbf{x}\cup\mathbf{y}$. A system ${X}$ shares a constraint or objective function with a system ${Y}$ if a variable property from both ${X}$ and ${Y}$ is included in the domain of that constraint or function. We call a subsystem fully consistent if all constraint and objective functions that are shared between ${X}$ and ${Y}$ are fully consistent.
V THEORETICAL RESULTS
In this section, we prove that optimizing a consistent subsystem and using its PF to optimize the larger system still yields globally optimal results, the basis for our monotone system decomposition technique. This accelerates computation as we can precompute subsystem solutions and leverage this reduced set to efficiently solve the larger design problem, drastically reducing the search space.
Theorem: Consider the optimization of a system ${Z}={X}\cup{Y}$ where ${Z}\subseteq{V}$ and ${X}$ is a fully consistent subsystem within ${Z}$. Then all assignments of ${Z}$ on the Pareto front $O_{Z}$ can use assignments from the Pareto front $O_{X}$ of the subsystem ${X}$, where $O_{X}$ is optimized for variable properties used in shared functions between ${X}$ and ${V}\setminus{X}$.
Subsystem consistency ensures that variable properties are optimized consistently, minimizing $f_{i}^{\downarrow}$ and $L_{c}$ and maximizing $f_{i}^{\uparrow}$ and $R_{c}$. This guarantees that, for any $\mathbf{x^{\prime}}\notin O_{X}$, there is an assignment $\mathbf{x}\in O_{X}$ that will have better objective values and will satisfy the same inequality constraints in $Z$.
Proof: As ${X}$ is a fully consistent subsystem, its shared properties are either maximization or minimization properties.
Consider the maximization property $D^{\uparrow}_{x,p}\in D_{x}$ in $X$. Let $\mathbf{x^{\prime}}$ and $\mathbf{x}$ be assignments such that $\mathbf{x^{\prime}}\notin O_{X}$, ${\mathbf{x}}\in O_{X}$. Thus, $d^{\uparrow}_{\mathbf{x^{\prime}},p}\leq d^{\uparrow}_{\mathbf{{x}},p}$ for $d^{\uparrow}_{\mathbf{x^{\prime}},p},d^{\uparrow}_{\mathbf{{x}},p}\in D^{\uparrow}_{x,p}$ as $O_{X}$ was maximized for $D^{\uparrow}_{x,p}$. Additionally, let $f^{\uparrow}_{i}(D_{{x},p},D_{y,p^{\prime}},\ldots)$ and $f_{i}^{\downarrow}(D_{{x},p},D_{y,p^{\prime}},\ldots)$ represent objective functions in the system ${Z}$, and let $R_{c}(D_{x,p},D_{y,p^{\prime}},...)$ and $L_{c}(D_{x,p},D_{y,p^{\prime}},...)$ represent shared constraint functions in ${Z}$.
We first prove optimality, in that an assignment $\mathbf{z}=\mathbf{x}\cup\mathbf{y},\mathbf{y}\in\mathbf{Y}$ will dominate an assignment $\mathbf{z^{\prime}}=\mathbf{x^{\prime}}\cup\mathbf{y}$. We then prove feasibility, in that a solution $\mathbf{z}$ will satisfy all constraints met by $\mathbf{{z^{\prime}}}$.
Optimality: Since the objective functions $f_{i}^{\uparrow}$ and $f_{i}^{\downarrow}$ are fully consistent and $D^{\uparrow}_{x,p}$ is a maximization property, it follows $f^{\uparrow}_{i}(d^{\uparrow}_{\mathbf{x^{\prime}},p},d_{\mathbf{y},p^{\prime}},\ldots)\leq f^{\uparrow}_{i}(d^{\uparrow}_{\mathbf{x},p},d_{\mathbf{y},p^{\prime}},\ldots)$ and $f^{\downarrow}_{i}(d^{\uparrow}_{\mathbf{x},p},d_{\mathbf{y},p^{\prime}},\ldots)\leq f^{\downarrow}_{i}(d^{\uparrow}_{\mathbf{x^{\prime}},p},d_{\mathbf{y},p^{\prime}},\ldots)$.
Feasibility: Since the constraints are structured $L_{c}(D_{x,p},D_{y,p^{\prime}},...)\leq R_{c}(D_{x,p},D_{y,p^{\prime}},...)$, then $\mathbf{x}$ will not violate any constraints that $\mathbf{x^{\prime}}$ does not already violate since $L_{c}(d^{\uparrow}_{\mathbf{x},p},d_{\mathbf{y},p^{\prime}},...)\leq L_{c}(d^{\uparrow}_{\mathbf{x^{\prime}},p},d_{\mathbf{y},p^{\prime}},...)$ and $R_{c}(d^{\uparrow}_{\mathbf{x^{\prime}},p},d_{\mathbf{y},p^{\prime}},...)\leq R_{c}(d^{\uparrow}_{\mathbf{x},p},d_{\mathbf{y},p^{\prime}},...)$ as $R_{c}$ and $L_{c}$ are fully consistent.
Thus, for a shared maximization property $D^{\uparrow}_{x,p}$, we have proven for any $\mathbf{z^{\prime}}=\mathbf{x^{\prime}}\cup\mathbf{y},\mathbf{x^{\prime}}\notin{O_{X}},\mathbf{y}\in{\mathbf{Y}}$, there is an assignment ${\mathbf{z}}=\mathbf{x}\cup\mathbf{y},\mathbf{x}\in{O_{X}}$ such that ${\mathbf{z}}\preceq\mathbf{{z^{\prime}}}$. The above holds for all $D^{\uparrow}_{x,p}$ shared between $X$ and $Y$. For minimization properties $D^{\downarrow}_{x,p}$, the proof is complementary but omitted due to space limitations. Since ${\mathbf{z}}\preceq\mathbf{{z^{\prime}}}\forall\mathbf{x}\in O_{X},\mathbf{x^{\prime}}\notin O_{X}$, and by definition $O_{Z}=\{\mathbf{z}\in\mathbf{Z}|\nexists\mathbf{z^{\prime}}\in\mathbf{Z},\mathbf{z^{\prime}}\preceq\mathbf{z}\}$, then ${O}_{Z}$ will use assignments from ${O_{X}}$. $\blacksquare$
VI DISCUSSION
In this section, we highlight advantages of monotone subsystem decomposition, address optimization of non-monotone systems, provide practical considerations, and compare our approach to MCDPs.
VI-A Advantages of Using Subsystems for Optimization
Using subsystems for optimization provides several advantages. First, subsystems are an intuitive way of describing the larger problem, and this framework is already commonly used by engineers to build complex systems.
Second, optimizing subsystems can provide a designer with internal information that can be helpful for system design. For example, in examining the PF of a sensing/computing subsystem, a designer may realize that a larger design problem is failing to meet design specifications since the subsystem does not have enough computing power.
Third, subsystems can be reused within the same or across different design problems, saving on computation. If the larger design problem changes or the subsystem is applied to a different design problem, the subsystem’s PF can be reused as long as its minimization and maximization properties remain unchanged. Furthermore, subsystems replicated in a larger design system (e.g. quadcopter designs used in a quadcopter delivery problem, see Sec. VII-B) can be optimized once and the resulting PF used as an “aggregate” component in the larger system, avoiding redundant computations.
VI-B Optimizing Non-Monotone Systems
Systems containing constraint or objective functions that are not fully consistent (e.g., non-monotone relationships or equality constraints) can still be optimized with this method after additional considerations.
First, we note that functions that are not fully consistent only violate the assumptions of the above proof if they are shared between the subsystem and the larger system. Inconsistent functions internal to the subsystem or exclusive to the larger system can still be optimized using standard CP methods without requiring consistency.
Otherwise, if these inconsistent functions are shared with the subsystem and assuming no other information on relationships, constraints, or properties, then every possible value of the inconsistent property must be evaluated. For example, if logic voltages appear in equality constraints, the system must be optimized for each logic voltage in the catalog. If the property is drawn from a relatively small finite set, the resulting PF may remain manageable. However, if the property spans a large finite set, possibly all subsystem implementations could be represented on the PF. Thus, selecting the proper subsystem is critical for tractability.
VI-C Practical Considerations for Using Subsystems
We provide guidelines for selecting effective subsystems, but in general this task depends on the details of a specific problem. First, an ideal subsystem has as few external properties as possible. Fewer external properties results in fewer optimization objectives, which is important for tractability as the solution space for many-objective problems scales rapidly. Second, it is easiest and most insightful for a designer to choose intuitive subsystems (e.g., the powertrain of a mobile robot) as these are more likely to be used in other systems (e.g., most robot designs would like to maximize the efficiency of a powertrain while minimizing cost). Finally, we note that even a single component can be made into a subsystem, which can be extremely beneficial as that component’s catalog can be optimized for a PF before other computation, potentially reducing the search space of the larger design problem.
VI-D Comparison of Our Approach to MCDPs
Our CP approach with monotone subsystem decomposition is similar to MCDPs in that they both use monotonicity and can solve multi-objective optimization problems; however, they differ in a few aspects. First, MCDPs can easily express continuous and discrete domains while our approach is limited to discrete domains but is better suited for the discrete problem of component selection [2]. Second, our monotone subsystem decomposition imposes fewer models restrictions. MCDPs require all relationships between variables to be monotone. Our approach allows both monotone and antitone relationships without requiring monotonicity throughout the entire system. In our approach’s framework, a discrete domain MCDP would be a consistent system with only maximization properties. Finally, CP offers the advantage of interfacing with other traditional CP problems, such as packing or scheduling (as demonstrated in the quadcopter fleet design problem in Sec. VII-B).
VII Computational Experiments
In this section, we first design a quadcopter system and compare our CP approach to a linear programming approach, the current state-of-the-art in terms of computational speed. We then expand upon this quadcopter design problem and solve a multi-objective, task-oriented scheduling problem to design a fleet of quadcopters to complete package deliveries as cheaply and as quick as possible.
In practice, due to the scale of our catalogs, there can exist several alternative solutions at a single point on the PF. We provide only one solution per point under the assumption that if a specific point interests a designer, then all solutions can be returned for that point with further processing.
All provided run times in this section were collected on a standard office desktop computer (4 GHz CPU with 12 cores, 32 GB RAM).
VII-A Quadcopter System Design Problem
We compare our CP approach to the LP method of Carlone et al. [3]. We use the same quadcopter design problem, a single-objective problem to design a quadcopter of maximum velocity within a fixed budget of \$1k. The quadcopter is composed of five types of components: motor, computer, camera, battery, and frame (see [3] for full model and constraints). We highlight that there are non-monotone constraints within the model. For example, velocity is a non-monotone quantity: the goal is to maximize velocity as that is the primary design objective, but within the constraints this quantity needs to be minimized since the camera’s minimum frame rate and the computer’s minimum back-end throughput are directly proportional to the velocity.
Carlone et al. use catalogs that have 19, 4, 4, 14, and 6 items for the motor, computer, camera, battery, and frame, respectively. To better assess scalability of each approach, we generate synthetic catalogs of 100,000 components per type based on the mean and standard deviation of the real-world components from the original catalogs. This yields problem sizes of up to 1025 component combinations.
Figure 3: Comparison of our CP approach versus the linear programming approach. At larger catalog sizes, the CP produces solutions faster than LP, and the multi-objective CP even produces an entire PF faster than the time it takes LP to determine one single-objective solution. Above 30,000 components, the LP application crashes.
Figure 3 compares the scaling performance of our CP approach against the LP approach. While the original quadcopter design problem is single-objective, our method extends to multi-objective optimization so we also report timing results for solving a multi-objective variant that minimizes system mass alongside maximizing velocity. As shown, our CP approach produces solutions faster than LP for all but the smallest of catalogs (in which solve times are less than a second). For catalogs exceeding 200 components, our CP approach even solves the multi-objective design problem and determines a PF of optimal solutions faster than the LP approach can solve the single-objective case. Our approach has a log-log slope $<1$, indicating its scaling is sub-linear.
VII-B Quadcopter Fleet Design Problem
We then applied our work to a larger and more complex case study solving the multi-objective, task-oriented scheduling problem of building a fleet of quadcopters and assigning them to deliver packages, all while optimizing for the delivery time and cost to build the fleet. We also solve this problem with a single-objective (only minimizing cost) to showcase the speed and scalability of our approach.
Figure 4: System diagrams for the quadcopter (a) and quadcopter fleet (b) design problems. The quadcopter contains a motor subsystem, which itself has an ESC subsystem. Several instantiations of the quadcopter serve as subsystems in the quadcopter fleet design problem, where each quadcopter is assigned packages for delivery.
This larger fleet design problem is broken up into several smaller design problems, see Figure 4. The highest level of abstraction entails the quadcopter and package assignment. Quadcopters are assigned to deliver packages. Packages all go to unique locations, and after delivery a quadcopter must return to the depot to pick up another package for delivery. A maximum number of unique quadcopter designs can be used; however, if more than one quadcopter design is chosen, an extra financial cost is incurred for each additional design.
The second level of abstraction is the quadcopter subsystem, nearly identical from Sec VII-A, with two adjustments. First, the quadcopter constraints formulated by [3] assume a racing quadcopter; since we are performing package delivery, we modify the camera and computer constraints to better reflect that these quadcopters will be built for landing at low speeds, not cornering at high speeds. Second, the motors of the original quadcopter subsystem are replaced with a new motor subsystem, the third level of abstraction. Carlone et al. assumed the motor, propeller, and electronic speed controller (ESC) were all integrated into a single “motor” component. Here, we break this out as a motor subsystem composed of these three components. The ESC itself is also a subsystem and the fourth level of abstraction. It is composed of an H-bridge, microcontroller, and voltage regulator. Each of these subsystems has several constraints that ensure component compatibility and adherence to target specifications.
First, to demonstrate the computational advantage of using subsystems, we recorded solve times for the quadcopter subsystem with and without subsystem decomposition. The quadcopter subsystem can be broken down into 11 subsystems: the ESC and motor subsystems, and one subsystem for each of the 9 components in the ESC, motor, and quadcopter subsystems. We used catalogs of 1,000, 250, and 1,000 components for each component in the ESC, motor, and quadcopter subsystems, respectively. With monotone subsystem decomposition, it takes 714.3 seconds to solve for the quadcopter subsystem PF, a 93% speed-up compared to the 1375.1 seconds to solve without subsystem decomposition.
Next, we separately optimized each subsystem in Fig. 4 and then used these results to optimize the larger design problem. Table I shows the number of items in the catalogs for each component type and the solve times for each subsystem. It takes 167.1 seconds to optimize for all of the subsystems, and these results can be subsequently reused in both the single- and multi-objective design problems to solve them in 1.3 and 352.9 seconds, respectively.
System Subsystem/ Component Number of Components System Solve Time
Electronic Robot Design Voltage Reg. 1000 43.4 sec
Speed H Bridge 1000 309 solns
Controller $\mu$-controller 1000
Motor ESC 309 102.7 sec
Propeller 250 179 solns
Motor 250
Quadcopter Motor Subsystem 179 21.0 sec
Battery 500 32 solns
Frame 500
Computer 500
Camera 500
Quadcopter Planning Quadcopter 25 1.3 sec
Fleet Packages 125 1 soln
Single-Obj. Designs 10 of 32
Quadcopter Quadcopter 4 352.9 sec
Fleet Packages 12 6 solns
Multi-Obj. Designs 3 of 32
TABLE I: System Components and Solve Times
VIII CONCLUSION
In this paper, we introduce a novel CP-based technique to optimize subsystems and leverage these PFs to efficiently solve larger multi-objective design problems, ensuring global optimality and providing useful design insights about internal subsystem design choices. Although this decomposition exploits monotonicity for computational gains, CP can still handle more general problems when monotonicity is absent. We demonstrate that our approach outperforms the previous state-of-the-art by an order of magnitude, and applied it to a larger, task-oriented scheduling problem optimizing a fleet of quadcopters for package delivery.
Currently, as the solver iteratively finds solutions on the PF, it must reset each time so that it can add a constraint that restricts that previous solution in subsequent solves. In future work, we intend to further speed up the solver by dynamically adding constraints as the solver is running. We also intend to investigate techniques of determining adjacent solutions once a solution has been found.

### v2

@@
 We index the variable properties $d_{v}$ using an index set ${P}_{v}=\{1,\ldots,N_{D_{v}}\}$, defined for each $v\in V$, so that the p-th property of a variable v can be written $D_{v,p}$.
-Constraints are mathematical expressions that describe relationships between the properties of components.
-These constraints can not only specify valid component combinations but also describe the performance specifications of the robot.
-Inequality constraints take the general form of $L_{c}(D_{v,p},D_{v^{\prime},p^{\prime}},...)\leq R_{c}(D_{v,p},D_{v^{\prime},p^{\prime}},...)$ where $L_{c}:\mathbb{R}^{\sum_{v}|P_{v}|}\rightarrow\mathbb{R}$ and $R_{c}:\mathbb{R}^{\sum_{v}|P_{v}|}\rightarrow\mathbb{R}$ are real-valued functions that operate on the variable properties and describe constraint $c\in C$.
+[…]
 The goal is to find an assignment $\mathbf{v}$ that maximizes or minimizes an objective function $f_{i}(\mathbf{v})$ subject to the constraints.
@@
 This shorthand allows constraints to be conveniently expressed as:
-$4*(M_{\textrm{voltage}}*M_{\textrm{current}})\vspace{-1.25mm}\leq B_{\textrm{voltage}}*B_{\textrm{current}}$
+$4*(M_{\textrm{voltage}}*M_{\textrm{current}})\vskip-3.55658pt\leq B_{\textrm{voltage}}*B_{\textrm{current}}$
 where the terms of the left and right hand side refer to the associated voltage and current properties of the motor (M) and battery (B) variables, respectively.
@@
 A constraint function $L_{c}$, $R_{c}$, or objective function $f_{i}$ is fully consistent if all variable properties in its domain are consistent.
-We can determine a subsystem ${X}$ within the constraint graph by selecting a subset of variables from the larger system ${Z}={X}\cup{Y}$, where ${X}\subset{Z}\subseteq{V}$ and ${Y}={Z}\setminus{X}$.
-We denote an implementation $\mathbf{z}$ of subsystem ${Z}$ that uses implementations $\mathbf{x}$ and $\mathbf{y}$ from ${X},{Y}$ as $\mathbf{z}=\mathbf{x}\cup\mathbf{y}$.
-A system ${X}$ shares a constraint or objective function with a system ${Y}$ if a variable property from both ${X}$ and ${Y}$ is included in the domain of that constraint or function.
-We call a subsystem fully consistent if all constraint and objective functions that are shared between ${X}$ and ${Y}$ are fully consistent.
+We can determine a subsystem X{X} within the constraint graph by selecting a subset of variables from the larger system ${Z}={X}\cup{Y}$, where ${X}\subset{Z}\subseteq{V}$ and ${Y}={Z}\setminus{X}$.
+We denote an implementation $\mathbf{z}$ of subsystem Z{Z} that uses implementations $\mathbf{x}$ and $\mathbf{y}$ from X,Y{X},{Y} as $\mathbf{z}=\mathbf{x}\cup\mathbf{y}$.
+A system X{X} shares a constraint or objective function with a system Y{Y} if a variable property from both X{X} and Y{Y} is included in the domain of that constraint or function.
+We call a subsystem fully consistent if all constraint and objective functions that are shared between X{X} and Y{Y} are fully consistent.
 V THEORETICAL RESULTS
@@
 This accelerates computation as we can precompute subsystem solutions and leverage this reduced set to efficiently solve the larger design problem, drastically reducing the search space.
-Theorem: Consider the optimization of a system ${Z}={X}\cup{Y}$ where ${Z}\subseteq{V}$ and ${X}$ is a fully consistent subsystem within ${Z}$.
-Then all assignments of ${Z}$ on the Pareto front $O_{Z}$ can use assignments from the Pareto front $O_{X}$ of the subsystem ${X}$, where $O_{X}$ is optimized for variable properties used in shared functions between ${X}$ and ${V}\setminus{X}$.
+Theorem: Consider the optimization of a system ${Z}={X}\cup{Y}$ where ${Z}\subseteq{V}$ and X{X} is a fully consistent subsystem within Z{Z}.
+Then all assignments of Z{Z} on the Pareto front $O_{Z}$ can use assignments from the Pareto front $O_{X}$ of the subsystem X{X}, where $O_{X}$ is optimized for variable properties used in shared functions between X{X} and ${V}\setminus{X}$.
 Subsystem consistency ensures that variable properties are optimized consistently, minimizing $f_{i}^{\downarrow}$ and $L_{c}$ and maximizing $f_{i}^{\uparrow}$ and $R_{c}$.
 This guarantees that, for any $\mathbf{x^{\prime}}\notin O_{X}$, there is an assignment $\mathbf{x}\in O_{X}$ that will have better objective values and will satisfy the same inequality constraints in $Z$.
-Proof: As ${X}$ is a fully consistent subsystem, its shared properties are either maximization or minimization properties.
-Consider the maximization property $D^{\uparrow}_{x,p}\in D_{x}$ in $X$.
-Let $\mathbf{x^{\prime}}$ and $\mathbf{x}$ be assignments such that $\mathbf{x^{\prime}}\notin O_{X}$, ${\mathbf{x}}\in O_{X}$.
-Thus, $d^{\uparrow}_{\mathbf{x^{\prime}},p}\leq d^{\uparrow}_{\mathbf{{x}},p}$ for $d^{\uparrow}_{\mathbf{x^{\prime}},p},d^{\uparrow}_{\mathbf{{x}},p}\in D^{\uparrow}_{x,p}$ as $O_{X}$ was maximized for $D^{\uparrow}_{x,p}$.
-Additionally, let $f^{\uparrow}_{i}(D_{{x},p},D_{y,p^{\prime}},\ldots)$ and $f_{i}^{\downarrow}(D_{{x},p},D_{y,p^{\prime}},\ldots)$ represent objective functions in the system ${Z}$, and let $R_{c}(D_{x,p},D_{y,p^{\prime}},...)$ and $L_{c}(D_{x,p},D_{y,p^{\prime}},...)$ represent shared constraint functions in ${Z}$.
+Proof: As X{X} is a fully consistent subsystem, its shared properties are either maximization or minimization properties.
+[…]
 We first prove optimality, in that an assignment $\mathbf{z}=\mathbf{x}\cup\mathbf{y},\mathbf{y}\in\mathbf{Y}$ will dominate an assignment $\mathbf{z^{\prime}}=\mathbf{x^{\prime}}\cup\mathbf{y}$.
@@
 Optimality: Since the objective functions $f_{i}^{\uparrow}$ and $f_{i}^{\downarrow}$ are fully consistent and $D^{\uparrow}_{x,p}$ is a maximization property, it follows $f^{\uparrow}_{i}(d^{\uparrow}_{\mathbf{x^{\prime}},p},d_{\mathbf{y},p^{\prime}},\ldots)\leq f^{\uparrow}_{i}(d^{\uparrow}_{\mathbf{x},p},d_{\mathbf{y},p^{\prime}},\ldots)$ and $f^{\downarrow}_{i}(d^{\uparrow}_{\mathbf{x},p},d_{\mathbf{y},p^{\prime}},\ldots)\leq f^{\downarrow}_{i}(d^{\uparrow}_{\mathbf{x^{\prime}},p},d_{\mathbf{y},p^{\prime}},\ldots)$.
-Feasibility: Since the constraints are structured $L_{c}(D_{x,p},D_{y,p^{\prime}},...)\leq R_{c}(D_{x,p},D_{y,p^{\prime}},...)$, then $\mathbf{x}$ will not violate any constraints that $\mathbf{x^{\prime}}$ does not already violate since $L_{c}(d^{\uparrow}_{\mathbf{x},p},d_{\mathbf{y},p^{\prime}},...)\leq L_{c}(d^{\uparrow}_{\mathbf{x^{\prime}},p},d_{\mathbf{y},p^{\prime}},...)$ and $R_{c}(d^{\uparrow}_{\mathbf{x^{\prime}},p},d_{\mathbf{y},p^{\prime}},...)\leq R_{c}(d^{\uparrow}_{\mathbf{x},p},d_{\mathbf{y},p^{\prime}},...)$ as $R_{c}$ and $L_{c}$ are fully consistent.
+[…]
 Thus, for a shared maximization property $D^{\uparrow}_{x,p}$, we have proven for any $\mathbf{z^{\prime}}=\mathbf{x^{\prime}}\cup\mathbf{y},\mathbf{x^{\prime}}\notin{O_{X}},\mathbf{y}\in{\mathbf{Y}}$, there is an assignment ${\mathbf{z}}=\mathbf{x}\cup\mathbf{y},\mathbf{x}\in{O_{X}}$ such that ${\mathbf{z}}\preceq\mathbf{{z^{\prime}}}$.
@@
 For catalogs exceeding 200 components, our CP approach even solves the multi-objective design problem and determines a PF of optimal solutions faster than the LP approach can solve the single-objective case.
-Our approach has a log-log slope $<1$, indicating its scaling is sub-linear.
+Our approach has a log-log slope <1<1, indicating its scaling is sub-linear.
 VII-B Quadcopter Fleet Design Problem